# INTENTAS collaboration INTENTAS - An entanglement-enhanced atomic sensor for microgravity

### v1

@@ IV.2 Vacuum system
The main vacuum for the system is provided by an ion-getter-pump (VACION PLUS 20 DIODE), which achieves a pumping speed of 27 l/s (N2). It is supported by a titanium sublimation pump (ZST22L110 VACGEN) which is irregularly refreshed. Before the first operation, the system is baked uniformly at 60 C and evacuated with a turbo molecular pump, which is removed afterwards from the system by pinching off the copper tube connections. The pressure can be monitored by a cold-cathode gauge (Agilent IMG-300) which is located directly at the support chamber but will be turned off during science operation to avoid spurious magnetic fields. The 3D-MOT is separated from the 2D+-MOT by a differential pumping section as used in [29]. This device of 63.5 mm length includes a graphite capillary which tapers from 8.8 mm to 1.5 mm over its extent and thereby reduces the pumping speed such that the pressure in the 2D+-MOT chamber is about three orders of magnitude higher compared to the science chamber. Additionally, due to the graphite inserts, an enhanced absorption is expected for Rb atoms, reducing the partial pressure of 87Rb in the science chamber further. The target pressure at the designated location of the atoms is below 10-8 Pa.
@@ IV.3 Magnetic shielding
The inner and outer layer of the walls are made out of twelve (inside: 404 mm 298 mm 2 mm, outside: 488 mm 364 mm 4 mm) mu-metal plates and provide shielding for static and low-frequency magnetic fields. The center structure, which is used to anchor the mu-metal plates, is made out of aluminum of size 446 mm 327 mm 5 mm, which provides further shielding against high-frequency (30 Hz - 100 MHz) magnetic fields. The magnetic shielding forms a volume of 0.176 m3 and with its additional top and bottom section encapsulates the entire sensor head. In order to provide an optical access for the detection system a tube of length 15 cm is added. In addition, six feedthroughs with the length of 19 cm were placed for optical fibers, cables and cooling lines, as well as a dedicated feedthrough for the ion-getter pump, which also acts as electrical ground for the science chamber. The mu-metal plates will be annealed before usage and equilibrated in all three dimensions using copper coils in an L-shape configuration around the inner and outer walls of the mu-metal in order to minimize build-up of residual magnetic fields [30]. The system has been simulated using ANSYS Academic 2022 for the given magnetic field in the Einstein-Elevator, which has been measured and evaluated in a test flight using a magnetometer (Bartington Mag-03) as shown in Fig. 4(a). The evaluation has revealed that the magnetic field strength change is less than 10 µT along the entire trajectory, while the largest offset field is around 70 µT. The result of the simulated reduction for the measured magnetic field in the EE can be seen in Fig. 4(b). The mean magnetic field strength according to the simulation is reduced by a factor for low-frequency fields (0.1 Hz) and is suppressed further with increasing frequency. This will yield a magnetic field strength of nT at the location of the atoms while operating in the Einstein-Elevator.
@@ V.1.1 Laser sources
Conceptually, the lasers are based on a master-oscillator-power-amplifier setup. In order to achieve a low frequency noise, an extended cavity diode laser (ECDL) serves as master oscillator. The operating parameters such as the injection current applied to the ECDL (I mA), the temperature set points of the micro-optical bench (T °C) and the volume holographic Bragg grating (T °C) are selected so that the laser emits in a stable single mode at the desired frequency (working point). Fig. 7(b) depicts the frequency noise of a science laser at the intended working point measured with a self-heterodyne setup. The white noise floor of the measurement [33] corresponds to an intrinsic linewidth of 2 kHz. Using the beta-separation line method [34] a FWHM-linewidth of 60 kHz (1 ms) can be derived.
After passing through a semi-double stage optical isolator, the emission of the ECDL is amplified in a semiconductor ridge-waveguide amplifier before it is coupled to a single-mode polarization maintaining optical fiber. Fig. 7(a) shows the optical power emitted from the output fiber for varying injection current into the amplifier at the intended working point of the laser. For operating the sensor in the EE an optical output power of 250 mW is required. Thus, an injection current into the ridge-waveguide amplifier of 750 mA was chosen as working point.
@@ V.1.2 Reference laser
As reference for the science lasers, a micro-integrated frequency reference module is used. It consists of a distributed feedback laser (DFB) with a dedicated frequency modulation spectroscopy (FMS) unit based on the design in [35, 36]. The design has been adapted to the needs of the sensor by modifying the laser and spectroscopy units and the included driver and control electronics. The module provides 15 mW of optical power via a polarization maintaining fiber and comes with a built-in laser current driver, temperature controller as well as lock electronics. A photograph of this subsystem can be seen in Fig. 8(a). The module is operated via a modified version of the Linien software [37]. It achieves a frequency stability of 1.7 10-12 after 1 s of integration time and 1.710-11 at a timescale of 8 h which was already shown in previous versions of the reference module [35, 36]. The stability of the module was measured over a time period of 13 h, by measuring the beat note with another lab based frequency reference. The module has a size of 135 mm 165 mm 40 mm and a mass of 3.3 kg.
@@ V.1.3 Beat detection system
The beat detection system is entirely fiber-based. The compartment housing this fiber network, is shown in Fig. 8(b). The fiber-based approach, with higher losses but also less need for maintenance, is preferred for long continuous measurement campaigns due to its expected stability gain in comparison with free-space optics. In the system, the light of the reference laser is split into four ports via an array of fiber splitters. The outputs of these four ports are mixed with the light of the science lasers in 50:50 fiber splitters. This generates the beat notes which are detected on fiber coupled photodiodes. The electrical signal is fed into the separated electronics compartment. The compartments are separated to reduce effects due to heating caused by the active electronic components. The module has a size of 210 mm 200 mm 87 mm and a mass of 1 kg.
@@ V.2.2 Magnetic field coils
| 2D+-MOT | 64 | 47x83 (rect.) | Quadrupole field | 56.5 µT/A/cm |
@@ V.3 Dipole trap
After the initial magneto-optical trapping and the subsequent polarization gradient cooling, the employed magnetic field and the cooling laser beams are turned off and the creation of the BEC will be achieved using a crossed optical dipole trap (cODT) in combination with time-averaged potentials.
@@ VI.1 Microwave source
The source providing the MW used for dressing and transitioning magnetic hyper-fine states is a copy of the system presented in [42]. The underlying schematic of the system is shown in Fig. 10. It provides two independently controllable output paths. Each path is able to provide 6.835 GHz with a low integrated phase noise of 480 µrad between 10 Hz and 100 kHz and an upper bound on the relative amplitude noise of 0.03. The system provides tunability in order to address different states depending on magnetic field of about 25 MHz. Additionally, it allows for fast dynamical adjustments in less than a µs. All of this is possible by combining a low phase noise oscillator with a configurable FPGA.
@@ VII.1 Experimental control
The control system is designed towards usage of the Advanced Real-Time Infrastructure for Quantum physics (ARTIQ) system which includes the Sinara hardware environment [43][44]. The hardware components of this system are variable and are only limited by spatial requirements of the capsule in the elevator. In general there is the possibility to use up to 3 FPGA modules (‘Kasli’), which allow several input/output operations via sub modules (e.g. DAC module ‘Zotino’). Furthermore, various RF frequency ramps of a few hundred MHz can be achieved using dedicated DDS devices (‘Urukul’ [45]) and higher frequencies (at least up to 12 GHz) are generated by the microwave synthesizer (‘Mirny’), which can be updated on the µs scale as well. Another module (‘Stabilizer’) acts independently as a PID control for the frequency stabilization and intensity stabilization of the lasers. All of this will be mounted into 19-inch rack spaces with custom made water cooling boards to dissipate the produced heat efficiently in the encapsulated environment. Next to the real-time system controlling the experimental sequences, there is a monitoring system based on the LabJack T7-Pro. This so-called slow control system is able to monitor relevant environmental parameters, e.g. temperature, and gives control over various hardware devices. The user interface for this relies on the Grafana environment [46].
@@ VIII Outlook
The INTENTAS apparatus offers flexible capabilities to enable a variety of experimental sequences and operations. The characteristic features, distinguishing it from other compact experiments, are the all-optical approach to BEC creation as well as the emphasis on providing all necessary tools to create an entangled atom source. The combination of the foreseen operation in microgravity and entanglement enhanced sensing opens up various applications.
@@
While there are several paths for entanglement possible, the further analysis is following the exemplary procedure in e.g. [24] to create spin-squeezed states. The squeezing parameter [47] can be estimated by solving the respective time-dependent Schrödinger equation using the Crank-Nicolson method. The result for 104 atoms is shown in Fig. 12. One can see that, within a reasonable range of trap frequencies, the time spin dynamics required to reach the best possible squeezing parameter does not limit the available time for the experimental sequence during the four seconds of free fall. Thus, one is able to take advantage of the microgravity environment in order to access extended Ramsey times in combination with entanglement-enhanced sensing. We estimate the possible enhancement in a prototypical interferometric measurement, assuming negligible technical noise. We choose a Ramsey measurement on the Rb clock transition as an example. While Ramsey spectroscopy on the Rb clock transition serves as a secondary definition of the SI second, it also forms the basis for Raman-based atom interferometers that enable the measurement of inertial moments. Fig. 13 shows the single-shot fractional frequency noise for three scenarios, without squeezing, with 10 dB squeezing that were demonstrated before [48], and with -21 dB as the theoretical optimum for our apparatus. Close to the optimal case (2 s Ramsey time and maximum squeezing) a single shot sensitivity of 1 10-14 would be accessible. However, even with more limited parameters the experiment can become competitive with state-of-the art compact atomic microwave clocks [49],[50]. This application demonstrates the advantages of harnessing the unique approach of an entangled compact source with extended Ramsey times in free fall. However, as Ramsey sequences on the Rb clock transition form a basis of both the operation of a microwave frequency standard and inertially sensitive interferometers based on Raman beam splitters, this is only an example and the extension towards the measurement of inertial moments by a Raman-based transfer to finite momentum states [51, 24] is foreseen.
IX Acknowledgements
@@
We would like to thank the QUANTUS and MAIUS teams for their insights and support. We are also thankful for the support of Waldemar Herr and Peter Fierlinger with regard to the magnetic shielding. We are also thankful for the support of Ioannis Papadakis with regard to the laser system. The INTENTAS-collaboration also would like to thank the Laboratory of Nano and Quantum Engineering (LNQE) for providing technology and the clean room environment for the assembly of the INTENTAS vacuum system. This work is supported by the German Space Agency (DLR) with funds provided by the BMWK under Grant No. 50WM2174, Grant No. 50WM2175, Grant No. 50WM2176, Grant No. 50WM2177, Grant No. 50WM2178. The authors would also like to thank the DFG and the Lower Saxony state government for their financial support for building the Hannover Institute of Technology (HITec) and the Einstein-Elevator (NI1450004, INST 187/624-1 FUGB) as well as the Institute for Satellite Geodesy and Inertial Sensing of the German Aerospace Center (DLR-SI) for the development and the provision of the carrier system. Supporting work was also contributed by the Deutsche Forschungsgemeinschaft (DFG) under Germany’s Excellence Strategy within the Cluster of Excellence QuantumFrontiers (EXC 2123, Project ID 390837967).

### v2

# INTENTAS - An entanglement-enhanced atomic sensor for microgravity

@@ I Introduction
The INTENTAS project paves the way for exploiting entanglement-enhanced interferometric sensitivities in a new generation of spaceborne atomic sensors.
@@ IV.2 Vacuum system
The main vacuum for the system is provided by an ion-getter-pump (VACION PLUS 20 DIODE), which achieves a pumping speed of 27 l/s (N2). It is supported by a titanium sublimation pump (ZST22L110 VACGEN) which is irregularly refreshed. Before the first operation, the system is baked uniformly at 60 °C and evacuated with a turbo molecular pump, which is removed afterwards from the system by pinching off the copper tube connections. The pressure can be monitored by a cold-cathode gauge (Agilent IMG-300) which is located directly at the support chamber but will be turned off during science operation to avoid spurious magnetic fields. The 3D-MOT is separated from the 2D+-MOT by a differential pumping section as used in [29]. This device of 63.5 mm length includes a graphite capillary which tapers from 8.8 mm to 1.5 mm over its extent and thereby reduces the pumping speed such that the pressure in the 2D+-MOT chamber is about three orders of magnitude higher compared to the science chamber. Additionally, due to the graphite inserts, an enhanced absorption is expected for Rb atoms, reducing the partial pressure of 87Rb in the science chamber further. The target pressure at the designated location of the atoms is below 10-8 Pa.
@@ IV.3 Magnetic shielding
In order to provide reproducible conditions in the experiment, any influence from changing magnetic fields have to be minimized. This is emphasized in an application where the external magnetic fields are changing due to the spatial movement of the experiment along different structures of the building. A passive magnetic shield is chosen to obtain an independence of the external conditions without the need of additional external coils for active compensation.
@@
The inner and outer layer of the walls are made out of twelve (inside: 404 mm 298 mm 2 mm, outside: 488 mm 364 mm 4 mm) mu-metal plates and provide shielding for static and low-frequency magnetic fields. The center structure, which is used to anchor the mu-metal plates, is made out of aluminum of size 446 mm 327 mm 5 mm, which provides further shielding against high-frequency (30 Hz - 100 MHz) magnetic fields. The magnetic shielding forms a volume of 0.176 m3 and with its additional top and bottom section encapsulates the entire sensor head. In order to provide an optical access for the detection system a tube of length 15 cm is added. In addition, six feedthroughs with the length of 19 cm were placed for optical fibers, cables and cooling lines, as well as a dedicated feedthrough for the ion-getter pump, which also acts as electrical ground for the science chamber. The mu-metal plates will be annealed before usage and equilibrated in all three dimensions using copper coils in an L-shape configuration around the inner and outer walls of the mu-metal in order to minimize build-up of residual magnetic fields [30]. The performance of the shield design has been simulated using ANSYS 2022. Therefore, the external magnetic field in the Einstein-Elevator has been measured and evaluated in a test flight using a magnetometer (Bartington Mag-03) as shown in Fig. 4. The evaluation has revealed that the magnetic field strength change is less than 5 µT along the entire trajectory, while the largest offset field is around 40 µT. Using those values, we find that the presented design of the magnetic shield leads to a substantial reduction for the measured magnetic field (see Fig. 5). Within the shield, at the location of the atoms, the mean magnetic field strength according to the simulation is expected to be reduced by a factor for low-frequency fields (0.1 Hz) and is suppressed further with increasing frequency. This will yield a magnetic field strength of nT at the location of the atoms while operating in the Einstein-Elevator.
@@ V.1.1 Laser sources
Conceptually, the lasers are based on a master-oscillator-power-amplifier setup. In order to achieve a low frequency noise, an extended cavity diode laser (ECDL) serves as master oscillator. The operating parameters such as the injection current applied to the ECDL (I mA), the temperature set points of the micro-optical bench (T °C) and the volume holographic Bragg grating (T °C) are selected so that the laser emits in a stable single mode at the desired frequency (working point). Fig. 9 depicts the frequency noise of a science laser at the intended working point measured with a self-heterodyne setup. The white noise floor of the measurement [33] corresponds to an intrinsic linewidth of 2 kHz. Using the beta-separation line method [34] a FWHM-linewidth of 60 kHz (1 ms) can be derived.
After passing through a semi-double stage optical isolator, the emission of the ECDL is amplified in a semiconductor ridge-waveguide amplifier before it is coupled to a single-mode polarization maintaining optical fiber. Fig. 8 shows the optical power emitted from the output fiber for varying injection current into the amplifier at the intended working point of the laser. For operating the sensor in the EE an optical output power of 250 mW is required. Thus, an injection current into the ridge-waveguide amplifier of 750 mA was chosen as working point.
@@ V.1.2 Reference laser
As reference for the science lasers, a micro-integrated frequency reference module is used. It consists of a distributed feedback laser (DFB) with a dedicated frequency modulation spectroscopy (FMS) unit based on the design in [35, 36]. The design has been adapted to the needs of the sensor by modifying the laser and spectroscopy units and the included driver and control electronics. The module provides 15 mW of optical power via a polarization maintaining fiber and comes with a built-in laser current driver, temperature controller as well as lock electronics. A photograph of this subsystem can be seen in Fig. 10. The module is operated via a modified version of the Linien software [37]. It achieves a frequency stability of 1.7 10-12 after 1 s of integration time and 1.710-11 at a timescale of 8 h which was already shown in previous versions of the reference module [35, 36]. The stability of the module was measured over a time period of 13 h, by measuring the beat note with another lab based frequency reference. The module has a size of 135 mm 165 mm 40 mm and a mass of 3.3 kg.
@@ V.1.3 Beat detection system
The beat detection system is entirely fiber-based. The compartment housing this fiber network, is shown in Fig. 11. The fiber-based approach, with higher losses but also less need for maintenance, is preferred for long continuous measurement campaigns due to its expected stability gain in comparison with free-space optics. In the system, the light of the reference laser is split into four ports via an array of fiber splitters. The outputs of these four ports are mixed with the light of the science lasers in 50:50 fiber splitters. This generates the beat notes which are detected on fiber coupled photodiodes. The electrical signal is fed into the separated electronics compartment. The compartments are separated to reduce effects due to heating caused by the active electronic components. The module has a size of 210 mm 200 mm 87 mm and a mass of 1 kg.
@@ V.2.2 Magnetic field coils
| 2D | 64 | 47x83 (rect.) | Quadrupole field | 56.5 µT/A/cm |
@@ VI.1 Microwave source
The source providing the MW used for dressing and transitioning magnetic hyperfine states is a copy of the system presented in [42]. The underlying schematic of the system is shown in Fig. 13. It provides two independently controllable output paths. Each path is able to provide 6.835 GHz with a low integrated phase noise of 480 µrad between 10 Hz and 100 kHz and an upper bound on the relative amplitude noise of 0.03. The system provides tunability in order to address different states depending on magnetic field of about 25 MHz. Additionally, it allows for fast dynamical adjustments in less than a µs. All of this is possible by combining a low phase noise oscillator with a configurable FPGA.
@@ VII.1 Experimental control
The control system is designed towards usage of the Advanced Real-Time Infrastructure for Quantum physics (ARTIQ) system which includes the Sinara hardware environment [43][44]. The hardware components of this system are variable and are only limited by spatial requirements of the capsule in the elevator. In general there is the possibility to use up to 3 FPGA modules (‘Kasli’), which allow several input/output operations via sub modules (e.g. DAC module ‘Zotino’). Furthermore, various RF frequency ramps of a few hundred MHz can be achieved using dedicated DDS devices (‘Urukul’ [45]) and higher frequencies (at least up to 12 GHz) are generated by the microwave synthesizer (‘Mirny’), which can be updated on the µs scale as well. Another module (‘Stabilizer’) acts independently as a PID control for the frequency stabilization and intensity stabilization of the lasers. All of this will be mounted into 19-inch rack spaces with custom made water cooling boards to dissipate the produced heat efficiently in the encapsulated environment.
Next to the real-time system controlling the experimental sequences, there is a monitoring system based on the LabJack T7-Pro. This so-called slow control system is able to monitor relevant environmental parameters, e.g. temperature, and gives control over various hardware devices. The user interface for this relies on the Grafana environment [46].
@@ VIII Outlook
The INTENTAS apparatus offers flexible capabilities to enable a variety of experimental sequences and operations. The characteristic features, distinguishing it from other compact experiments, are the all-optical approach to BEC creation as well as the emphasis on providing all necessary tools to create an entangled atom source. The combination of the foreseen operation in microgravity and entanglement-enhanced sensing opens up various applications.
VIII.1 Entanglement generation
The generation of entanglement can be achieved via different schemes and paths. One procedure to create spin-squeezed states is described in[24] and will be used as an example to demonstrate the capabilities of the setup. Here, the entanglement is created using atoms that are initialized in . By using microwave dressing, spin-changing collisions are induced transferring atoms into . Spin-changing collisions generate a two-mode-squeezed vacuum state in the levels . Afterwards, the atoms from are transferred to . Subsequently, a -polarized rf pulse transfers a single-mode squeezed substate from to . Together, the single-mode squeezed vacuum state in level and the large number of atoms in level form a two-mode spin-squeezed state. The remaining atoms in level are not coupled to the interferometer sequence and can be treated as a negligible loss channel. A tomography of the spin-squeezed state can be performed by a mw pulse (/2) and a subsequent measurement of the population imbalance. A variation of the mw phase enables a determination of the squeezing angle and the optimal squeezing parameter [47]. In order to quantify the theoretical limitations of our setup, one can estimate the parameter by solving the respective time-dependent Schrödinger equation using the Crank-Nicolson method [47]. The result for 104 atoms is shown in Fig. 15. One can see that, within a reasonable range of trap frequencies, the time that spin dynamics requires to reach the best possible squeezing parameter does not limit the available time for the experimental sequence during the four seconds of free fall. Thus, one is able to take advantage of the microgravity environment in order to access extended Ramsey times in combination with entanglement-enhanced sensing. We estimate the possible enhancement in a prototypical interferometric measurement, assuming negligible technical noise. Within the exemplary preparation described above, we choose a Ramsey measurement on the Rb clock transition as an example. In order to perform this, two pulses with the Ramsey time as delay between them are used after the entangled states are prepared.
VIII.2 Exemplary application: Ramsey spectroscopy
Ramsey spectroscopy on the Rb clock transition not only serves as a secondary definition of the SI second, it also forms the basis for Raman-based atom interferometers that enable the measurement of inertial moments. Therefore, we exemplify the targeted entanglement enhancement by a Ramsey spectroscopy sequence. In the experimental sequence of Fig. 14, we assume the generation of a BEC in level . The preparation is optionally followed by the generation of squeezing during the ”spin dynamics” phase and its transfer to the clock states during the ”state manipulation” phase. During the ”interrogation time”, we consider the application of two pulses on the Rb clock transition with a variable Ramsey time in between. A final measurement of the number of atoms in the two hyperfine states enables a referencing of the microwave frequency to the atomic hyperfine transition, which forms the basis for a microwave frequency standard.
@@ VIII.2 Exemplary application: Ramsey spectroscopy
Fig. 16 shows the single-shot fractional frequency noise for three scenarios, without squeezing, with 10 dB squeezing that were demonstrated before [48], and with -21 dB as the theoretical optimum for our apparatus. Close to the optimal case (2 s Ramsey time and maximum squeezing), a single-shot sensitivity of 1 10-14 would be accessible. However, even with more limited parameters, the experiment can become competitive with state-of-the art compact atomic microwave clocks [49],[50]. This application demonstrates the advantages of harnessing the unique approach of an entangled compact source with extended Ramsey times in free fall. However, as Ramsey sequences on the Rb clock transition form a basis of both the operation of a microwave frequency standard and inertially sensitive interferometers based on Raman beam splitters, this is only an example and the extension towards the measurement of inertial moments by a Raman-based transfer to finite momentum states [51, 24] is foreseen.
List of Abbreviations
| Abbreviation | Full Term |
| AOD | Acousto-Optical Deflector |
| AOM | Acousto-Optical Modulator |
| AR | Anti-Reflective |
| BEC | Bose-Einstein Condensate |
| DAC | Digital-to-Analog Converter |
| DDS | Direct Digital Synthesis |
| ECDL | External Cavity Diode Laser |
| EE | Einstein-Elevator |
| FPGA | Field-Programmable Gate Array |
| MOT | Magneto-Optical Trap |
| MW | Microwave |
| PER | Polarization Extinction Ratio |
| PID | Proportional-Integral-Derivative |
| RF | Radio Frequency |
| SWaP | Size, Weight, and Power |
IX Declarations
Ethics approval and consent to participate
Not applicable
Consent for publication
Not applicable
Availability of data and materials
Data are available upon reasonable request.
Competing interests
All authors declare that they have no competing interests.
Authors contributions
A.F. and J.S.H. drafted and edited the manuscript. O.A., A.F., M.G, J.H., J.S.H., A.H., C.K, M.S. contributed to the writing. C.K., J.S.H., J.H., A.F. J.K., E.M.R. contributed to the overall concept and design of INTENTAS. O.A., I.B., M.F., M.G., A.H., S.K, M.K, S.K, C.L, K.M., J.P., M.S. S.S., A.W. contributed to designs of subsystems. A.F., C.K., E.G., E.R., W.S., L.W. developed applications scenarios to deduce design requirements. All authors reviewed the manuscript.
Funding
@@ IX Declarations
This work is supported by the German Space Agency (DLR) with funds provided by the BMWK under Grant No. 50WM2174, Grant No. 50WM2175, Grant No. 50WM2176, Grant No. 50WM2177, Grant No. 50WM2178. The authors would also like to thank the DFG and the Lower Saxony state government for their financial support for building the Hannover Institute of Technology (HITec) and the Einstein-Elevator (NI1450004, INST 187/624-1 FUGB) as well as the Institute for Satellite Geodesy and Inertial Sensing of the German Aerospace Center (DLR-SI) for the development and the provision of the carrier system. Supporting work was also contributed by the Deutsche Forschungsgemeinschaft (DFG) under Germany’s Excellence Strategy within the Cluster of Excellence QuantumFrontiers (EXC 2123, Project ID 390837967).
Acknowledgements
We would like to thank the QUANTUS and MAIUS teams for their insights and support. We are also thankful for the support of Waldemar Herr and Peter Fierlinger with regard to the magnetic shielding. We are also thankful for the support of Ioannis Papadakis with regard to the laser system. The INTENTAS-collaboration also would like to thank the Laboratory of Nano and Quantum Engineering (LNQE) for providing technology and the clean room environment for the assembly of the INTENTAS vacuum system.